An object have mass, m is dropped from rest from a height, h. As it strikes the ground, it imparts an impulse to the ground proportional to( by assuming no energy loss)
A. m√h
B. mh
C. mh^2
D. None of above
1
Expert's answer
2020-09-17T14:29:09-0400

For our task "h=\frac{gt^2}{2}" , "v=gt" hence "v=\sqrt{2gh}" then "p=2m\sqrt{2gh}" where "p" is imparts an impuls to the ground.

Then "p" is proportional to "m\sqrt{h}" .